TREATING THE EXCEPTIONAL CASES OF THE MEATAXE

GÁBOR IVANYOS AND KLAUS LUX

Abstract. We show that the Holt–Rees extension of the standard MeatAxe procedure finds submodules of modules over finite algebras with positive probability in more cases than originally claimed. For the case when the Holt–Rees method fails we propose a further, but still simple and efficient extension.